OpenStudy (anonymous):

What is the sum of the geometric sequence 1, -4, 16, if there are 7 terms?

OpenStudy (amistre64):

find the next 4 terms and add them all up

OpenStudy (amistre64):

or, recall the formula for a geometric series ....

OpenStudy (anonymous):

The problem is I Dont understand how to do this, can u explain it to me ?

OpenStudy (amistre64):

well, a geometric series is one in which each new term is created by multipying the last one by some constant value ... 1, times what, is -4?

OpenStudy (anonymous):

-4

OpenStudy (amistre64):

good; and -4, times -4 is 16 so this creates the sequence we are looking for 1 -4 16 16*(-4) 16*(-4)*(-4) 16*(-4)*(-4)*(-4) 16*(-4)*(-4)*(-4)*(-4) those will be the first 7 terms, add them up

OpenStudy (amistre64):

otherwise, you would have to remember yet another formula ... 1 - r^n a * -------; a = first term, r = common ration, n=number of terms 1 - r

OpenStudy (anonymous):

Im Trying to add them up

OpenStudy (anonymous):

It gives me 3405

OpenStudy (amistre64):

close ... 3277 should be it

OpenStudy (anonymous):

Aw Okay , Thankx

OpenStudy (amistre64):

spose we have a geometric sequence: a, ar, ar^2, ar^3, ..., ar^n the sum of this sequence is: S= ar^0 + ar^1 + ar^2 + ar^3 + ... + ar^(n-1) multiply by r again we get: rS = ar^1 + ar^2 + ar^3 + ar^4 + ... + ar^n subtracting these we get: S = ar^0 + ar + ar^2 + ar^3 + ... + ar^(n-1) -rS = - ar - ar^2 - ar^3 - ... - ar^(n-1) - ar^n ----------------------------------------------- (1-r)S = a + 0 + 0 + 0 + ... + 0 - ar^n divide boths sides by 1-r S = (a - ar^n))/(1-r) factor out the a to get: S = a (1-r^n)/(1-r) for the formula

OpenStudy (anonymous):

Oh

OpenStudy (amistre64):

:) its fun to remember the math behind the formula ... just in case you forget the formula itself